OpenStudy (anonymous):

Solve for W in P=2W + 2Q, if P=38 and Q=12. W= ?

OpenStudy (anonymous):

substitute the values already given: 38= 2w +2(12) 38 = 2w + 24 then subtract: 38 = 2w + 24 -24 -24 14 = 2w then divide both sides by 2: 14/2=2w/2 7= w

OpenStudy (anonymous):

Okay thanks!